OpenStudy (anonymous):

6(3 - a) < 8a - 10 line under inequality sign

OpenStudy (anonymous):

expand 18-6a=<8a-10 add 10 28-6a=<8a add 6a 28=<14a divide by 14 2=<a

OpenStudy (saifoo.khan):

\[6(3 - a) \le 8a - 10\] \[18 - 6a \le 8a - 10\] \[8a-10 \ge 18 -6a\] \[8a + 6a \ge 18 + 10\] \[14a \ge 28\] \[a \ge2\]